4 years hence, the father age will be 3 times son's age. 6 years ago he was 8 times older than son. What is their present age?
Answer by lwsshak3(11628)

4 years hence, the father age will be 3 times son's age. 6 years ago he was 8 times older than son. What is their present age?
***
6 yrs ago:
x=son's age
8x=father's age
..
At present:
x+6=son's age
8x+6=father's age
..
4 yrs hence:
x+6+4=x+10=son's age
8x+6+4=8x+10=father's age
8x+10=3(x+10)
8x+10=3x+30
5x=20
x=4
Present ages:
x+6=10(son)
8x+6=38(father)